OpenStudy (anonymous):

The table shows values for points on the graph of a linear function. x: −2 1 2 4 y: 13 1 −3 −11 What is the slope of the graph of this linear function?

563blackghost (563blackghost):

Use the slope formula.... \(\LARGE{\frac{y2-y1}{x2-x1}}\) in which you would input two points.... \(\Huge\color{red}{\frac{1-13}{1-(-2)}=slope}\)

OpenStudy (triciaal):